# Zero modes and geometric phase for 2D Weyl fermions on Lifshitz backgrounds

### v1

@@ I Introduction
One can use semi-holographic [gursoy2013holographic] and holographic [landsteiner2016quantum, landsteiner2020holographic] methods to explore properties of Weyl semimetals from gravity descriptions in Lifshitz spacetimes. Indeed, in these setups, asymptotically Lifshitz black branes in $4+1$ dimensions are used to provide gravity descriptions of strongly coupled fermions on its respective $3+1$ dimensional boundaries. The dual field theory which emerges is nonrelativistic due to the inherent anisotropic scaling symmetry and, by addressing the problem to study Weyl fermions, we find a scenario where the boundary of the system behaves similarly to Weyl semimetals, since in these materials the Fermi velocity is small compared to the speed of light and in the vicinity of the Weyl nodes the emerging quasiparticles are effective massless fermions with well defined chirality. From this perspective, phenomena such as anomalous transport, chiral magnetic effects, and topological phase transitions have been analyzed from a holographic viewpoint [ammon2015gauge, Ahn2024_Praper, palumbo2016holographic]. Additionally, the recent experimental realization of two-dimensional Weyl semimetals [lu2024realization] brings to light the need to better understand the behavior of massless fermions on curved two-dimensional surfaces, as well as their links with different physical systems. Thinking along these lines, the holographic approach can provide insights into the physics of these new materials, helping to better understand them and opening up new branches for investigation.
In this work we are interested in studying some properties of Weyl fermions on $2+1$-dimensional Lifshitz backgrounds. These surfaces can be understood as boundaries of some extremal (zero temperature) 3+1 black-brane model equipped with scaling anisotropy. In these setups, we have strongly coupled nonrelativistic regimes, which implies that when addressing the study of massless fermions, we are led to systems which mimic some aspects of the behavior of 2D Weyl semimetals. In particular, we show that on these setups the Weyl spinor acquires a geometric (Berry) phase and presents a well-behaved zero mode, derived from an underlying supersymmetric quantum mechanics.
@@ III Holonomy and Wilson Loop
Note that $U(\mathcal{C})$ depends on the dynamical exponent $z$ and the scale factor $\ell$. Its associated topological invariant, the Wilson loop [cai1991neutrino, de2003loop], is related to the polarization of the Weyl fermion and can be expressed as
@@ IV Fermion zero mode
with $\alpha=1-1/z$ and $\tilde{\kappa}=kz^{-\alpha}$. The pair of equations (20a,b) can be factorized as
@@
which is normalizable for $\alpha<1$ (i.e., $z>0$), with $c_{1}=0$ if $\tilde{\kappa}>0$ or $c_{2}=0$ if $\tilde{\kappa}<0$. The existence of well-behaved zero modes indicates that the defects associated with the curvature of Lifshitz spacetime allow the emergence of zero-energy fermionic states in their vicinity.
@@ V Ending comments
This is an initial study concerning fermions on geometries with anisotropic scaling invariance, and we hope that further investigations will be useful for future works on effective gravitational models and their applications in planar condensed matter systems, since there are several connections between the physics of gravitational systems and the physics of materials such as graphene, topological insulators, and 2D Weyl semimetals. Within this context, several other phenomena can be explored, such as Landau levels, Hall effect and the Aharonov-Bohm effect, for example. We hope that these studies will open up a range of new research opportunities and be useful for a better understanding of the physics of two-dimensional materials.

### v2

@@ I Introduction
One can use semi-holographic [gursoy2013holographic] and holographic [landsteiner2016quantum, landsteiner2020holographic] methods to explore properties of Weyl semimetals from gravity descriptions in Lifshitz spacetimes. Indeed, in these setups, asymptotically Lifshitz black branes in $4+1$ dimensions are used to provide gravity descriptions of strongly coupled fermions on its respective $3+1$ dimensional boundaries. The dual field theory which emerges is nonrelativistic due to the inherent anisotropic scaling symmetry and, by addressing the problem to study Weyl fermions, we find a scenario where the boundary of the system behaves similarly to Weyl semimetals, since in these materials the Fermi velocity is small compared to the speed of light and in the vicinity of the Weyl nodes the emerging quasiparticles are effective massless fermions with well defined chirality. From this perspective, phenomena such as anomalous transport, chiral magnetic effects, and topological phase transitions have been analyzed from a holographic viewpoint [ammon2015gauge, Ahn2024_Praper, palumbo2016holographic]. Additionally, the recent experimental realization of two-dimensional Weyl semimetals [lu2024realization] highlights the need to expand our understanding on the behavior of massless fermions on curved two-dimensional surfaces, as well as their links with different physical systems. These surfaces can be understood as the boundary of an extremal 3+1 black brane model equipped with scaling anisotropy and studying the behavior of fields on these surfaces can give us more hints about the physics of strongly coupled non-relativistic quantum systems.
The discussion above indicates the existence of a close relation between the physics of Weyl fermions and the properties of Lifshitz spacetimes in the description of certain condensed matter phenomena. Therefore, thinking along these lines, one way to better understand this relationship is through the study of massless fermions on background geometries which exhibit anisotropic scaling isometries. Such analyses can provide new hints about the physics of these new materials, helping to better understand them and opening new avenues of investigation. Here we are interested in studying some properties of Weyl fermions on $2+1$-dimensional Lifshitz backgrounds. In particular, we show that on these setups the Weyl spinor acquires a geometric phase and presents a well-behaved zero mode, derived from an underlying supersymmetric quantum mechanics. Holonomy calculus studies the transport of a given quantity to the spin connection. This quantity depends on the path geometry and, for closed and adiabatic paths, it is usually called Berry geometric phase. But in a spacetime, we have several possibilities of closed paths, which involve all coordinates - time included. When one consider a given path involving a displacement in the time coordinate, a non-trivial holonomy associated with temporal translations is captured. This phase arises from the generalization on holonomy calculus, i.e., it is a generalization of the Berry phase [bezerra1992, gomes2007loop, de2003loop, de2025adiabaticity, cai1991neutrino]. The physical interpretation of this temporal holonomy is associated with a shift in energy due to time boosts associated to chemical potential and the emerging squeezed quantum states.
@@ III Holonomy and Wilson Loop
Note that $U(\mathcal{C})$ depends on the dynamical exponent $z$ and the scale factor $\ell$. Its associated topological invariant, the Wilson loop [cai1991neutrino, de2003loop, bezerra1992, gomes2007loop, de2025adiabaticity], is related to the polarization of the Weyl fermion and can be expressed as
@@ IV Fermion zero mode
with $\alpha=1-1/z$ and $\tilde{\kappa}=kz^{-\alpha}$. Note that since $z>1$, the values of the $\alpha$-parameter in the geometry-induced potential (21) are confined to the range $0<\alpha<1$. In particular, when approaching relativistic regimes considering the limit $z\to 1$, we have $\alpha\to 0$ and $V_{\pm}(r)\to\kappa^{2}=const.$, indicating that the reduction of the anisotropic scaling between the space and time coordinates implies the elimination of fermion bound states. In extremely anisotropic setups, where $\alpha\to 1$ for $z\to\infty$, we also find scenarios with no bound states, since in these cases we have $V_{\pm}(r)\to 0$. Therefore, one can observe that the presence of well-behaved fermion bound states requires the existence of anisotropy between space and time. However, if this anisotropy is too intense, such states may not survive. This may indicate that the electronic properties of 2D Weyl semimetals can change if their associated Fermi velocity is too low. In particular, for $z=2$, we have $\alpha=1/2$ and we find a scenario with a Coulomb potential complemented by a higher-order term.
The pair of equations (20a,b) can be factorized as
@@
which is normalizable since $\alpha<1$, with $c_{1}=0$ if $\tilde{\kappa}>0$ or $c_{2}=0$ if $\tilde{\kappa}<0$. The existence of well-behaved zero modes indicates that the defects associated with the curvature of Lifshitz spacetime allow the emergence of zero-energy fermionic states in their vicinity.
@@ V Ending comments
This is an initial study concerning fermions on geometries with anisotropic scaling invariance, and we hope that further investigations will be useful for future works on effective gravitational models and their applications in planar condensed matter systems, since there are several connections between the physics of gravitational systems and the physics of materials such as graphene, topological insulators, and 2D Weyl semimetals. Within this context, several other phenomena can be explored, such as Landau levels, Hall effect and the Aharonov-Bohm-like effect, for example. We hope that these studies will open up a range of new research opportunities and be useful for a better understanding of the physics of two-dimensional materials.